Using the formula PV = nRT and the following given information:

P = 0.520 atm (H2)

V = 30.0 mL

R = 0.69

T = 22.1 degrees Celsius

find the number of Hydrogen moles produced
Answer #1

Given:

P = 0.52 atm

V = 30.0 mL

= (30.0/1000) L

= 0.03 L

T = 22.1 oC

= (22.1+273) K

= 295.1 K

find number of moles using:

P * V = n*R*T

0.52 atm * 0.03 L = n * 0.08206 atm.L/mol.K * 295.1 K

n = 6.442*10^-4 mol

Answer: 6.44*10^-4 mol
Answer #2

To find the number of moles of hydrogen gas produced, we can rearrange the ideal gas law equation (PV = nRT) to solve for n (number of moles):

n = PV / RT

Given: P = 0.520 atm V = 30.0 mL = 0.0300 L (convert mL to L) R = 0.69 L·atm/(mol·K) T = 22.1 °C = 295.25 K (convert Celsius to Kelvin)

Substituting the values into the equation:

n = (0.520 atm * 0.0300 L) / (0.69 L·atm/(mol·K) * 295.25 K)

n ≈ 0.0079 moles

Therefore, approximately 0.0079 moles of hydrogen gas would be produced.


answered by: mervetokaz